Starting at the end of this year, you plan to make annual deposits of $8,000 for the next 10 years followed by deposits of $9,000 for the following 10 years. The deposits earn interest of 7.0%. What will the account balance be by the end of 25 years? Round to the nearest cent.
Answer #1
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
FV= 8000*(((1+ 7/100)^10-1)/(7/100))
FV = 110531.58
Future value = present value*(1+ rate)^time
Future value = 110531.38*(1+0.07)^15
Future value = 304959.56
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
FV= 9000*(((1+ 7/100)^10-1)/(7/100))
FV = 124348.03
Future value = present value*(1+ rate)^time
Future value = 124348.03*(1+0.07)^5
Future value = 174404.54

Total = 174404.54+304959.56=479364.1